OpenStudy (anonymous):

Hours worked - 19; Hourly rate - $5.75; What is the weekly pay?

OpenStudy (lieutenantgeneral):

19 * 7 = 133 $5.75 * 133 = $764.75 or $5.75 * 7 = $40.25

OpenStudy (anonymous):

Hours worked - 40; Hourly rate - $6.45; What is the weekly pay?

OpenStudy (anonymous):

Hours worked - 27; Total Pay - $125.55; What is the hourly rate? $4.65 $5.75 $4.75 26. Hours worked - 32; Total pay - $160; What is the hourly rate? $5.65 $4.75 $5.00 27. Hourly rate - $6.00; What is the overtime rate? $3.00 $9.00 $12.00 28. Hourly rate - $4.75; What is the overtime rate? $7.13 $9.50 $2.38 29. Aldo works for a moving company after school for 4 hours each day on Monday through Thursday afternoons. On Saturday, he works for 6 hours. At $7.50 per hour, with no overtime, how much does he earn each week? $195 $165 $185 30. On Saturdays, Mike earns $6.75 per hour working for 7 hours at an antique store. His boss offered him a $50 bonus and his regular rate of pay, if he worked for 9 hours on Sunday. How much did Mike earn for the weekend? $97.25 $147.25 $158.00 31. Peg works 2 1/2 hours each day on 4 days a week sewing covers. At $11.25 per hour, how much does she earn each week? $73.13 $90.00 $112.50 32. Kyle works in the stock department of a discount toy store for 8 hours each on Saturdays and Sundays. One weekend he earned $76. How much did he earn per hour? $4.75 $5.25 $5.75 33. Gershon is a machinist. He worked 45 hours in one week, of which 5 hours were overtime at time-and-a-half. He earned $463.60. How much did he earn per hour? $10.30 $9.76 $10.17 Part 7 Choose the correct earnings for each of the following exercises. (Each question is worth one point) 34. 102 units at $8.72 each $789.44 $919.44 $889.44 35. 2,019 units at $0.61 each $1,231.59 $11,315.90 $1,331.59 36. 865 units at $2.03 each $1730.55 $1655.95 $1755.95 37. $2.95 per hour; 30 hours; $295.25 tips $88.50 $206.75 $383.75 38. $1.95 per hour; 40 hours; $149.50 tips $78.00 $227.50 $149.50 39. $2.75 per hour; 35 hours; $207.40 tips $303.65 $96.25 $207.40 Part 8 Find the average weekly earnings for the following exercises, then choose the correct answer for each. (Each question is worth one point) 40. Hours: 29; Hourly rate: $4.33; Weekly tips: $135.98, $167.95, $198.25, $170.35 $293.52 $295.57 $293.70 41. Hours: 20; Hourly rate: $3.25; Weekly tips: $115.75, $94.80, $73.15, $62.50 $138.15 $151.55 $148.00 42. Unit rate: $0.76; Units per week: 502, 398, 433, 480 $344.47 $329.08 $364.80 43. Unit rate: $1.15; Units per week: 475, 348, 516, 402 $546.25 $462.30 $500.54 Part 9

OpenStudy (lieutenantgeneral):

OK.